OpenStudy (anonymous):

Use the Pythagorean Theorem to find the distance to the nearest tenth, between F(9, 5) and G(-2, 2). (Hint: Place F and G on the coordinate plane. Then sketch a triangle with FG as the hypotenuse.)

OpenStudy (anonymous):

Put a point H on (9, 2). Sketch a triangle out of the three points. Distance between (-2, 2) and (9, 2) is going to be 11. Distance between (9, 2) and (9, 5) is going to be 3. These correspond to the a and b of the Pythagorean Theorem \[c^2=a^2+b^2\] \[c = \sqrt{11^2+3^2}=\sqrt{130}\]

OpenStudy (anonymous):

Square root of 130 is 11.4

OpenStudy (anonymous):

well pythagorean theorem is a^2+b^2=c^2 so a is 11 because you do the distance between 9 and -2 then b is the distance between 5 and 2 which is 3 so you get 11^2 = 121 and 3^2 = 9 so sqrt of 130 is the answer which is between 11 and 12

OpenStudy (anonymous):

thanks!